OpenStudy (anonymous):

If the radius of a circle is doubled in length, the circumference of the circle is A. halved B. doubled C. tripled

OpenStudy (perl):

If you start with C = 2 π r Then you double the radius. C ' = 2 π (2r ) what happens to the new circumference C '

OpenStudy (perl):

C ' = 2 π (2r ) = 2 * 2π r = 2 * C

OpenStudy (anonymous):

thankyou @perl